# Anti-Phishing Training Does Not Work: A Large-Scale Empirical Assessment of Multi-Modal Training Grounded in the NIST Phish Scale

### v1

@@ 3 Methodology
In this section, we describe the methodology of our measurement study, including our hypotheses, experimental design, participant demographics, training modalities, and phishing simulations. An illustration is given in Figure 1. Our two-factor study design controls for the effect of training modality and task complexity. We make the first real-world application of the NIST Phish Scale framework [33], allowing us to evaluate the effectiveness of different training modalities across standardized difficulty levels in an operational environment.
TABLE III: Summary of study hypotheses, their theoretical grounding, and relation to prior work. The hypotheses address the effects of phishing email difficulty, training, and training modality on detection and reporting behaviors.
Hypothesis Theoretical and Empirical Basis Relation to Prior Work
H1 (Phish Scale Effect): Phishing emails with higher difficulty, as measured by the NIST Phish Scale, will exhibit higher click-through rates than those with lower difficulty. Grounded in the foundational NIST Phish Scale framework [33] and validation studies [39]. Supported by phishing cue research [34] and systematic email analysis [40]. The NIST Phish Scale has been validated in lab settings [39]. We provide the first large-scale enterprise validation.
H2 (Training Impact — General): Training will increase the detection and reporting of phishing emails. Evidence of training efficacy, reduced click rates [22], improved detection accuracy [7], and short-term benefits [21]. Disputed by a recent study [4]. Tests whether training effects from small-scale studies [22, 7] replicate at scale, addressing the call for large-sample validation [10].
H3 (Interactive Training Effect — Reporting): Interactive components (Trained+Exercise) will lead to significantly higher reporting rates than lecture-only training (Treatment A). From Franz et al.’s modality taxonomy [10], and evidence of the effectiveness of active and feedback-rich training [22, 21, 8]. Builds on academic findings of interactive training benefits [22, 21]. We provide the first test at organizational scale.
H4 (Interaction Effect): There will be a significant interaction between training modality and phishing difficulty, with interactive training being more beneficial for easier lures. Emerges from cognitive load and decision strategy research [19], difficulty-aware detection patterns [40], and findings on human limitations [6]. Addresses the gap identified by Franz et al. [10] regarding interaction effects between training and phishing complexity, largely unexplored empirically.
3.1 Hypotheses
Based on prior work on phishing training efficacy [22, 23, 5, 7, 10, 4] and documented differences in risk perception [6, 9], we formulated four hypotheses to test in this work. Table III states each hypothesis, its basis, and its relation to prior work.
To summarize, H1 represents the foundational assertion of the NIST Phish Scale framework [33]; H2 and H3 capture the intended effects of training interventions, with H3 specifically testing whether interactive modalities provide measurable benefits over passive learning approaches; and H4 reflects our expectation that training effectiveness may diminish for highly sophisticated phishing attempts, as advanced phishing emails may exceed individuals’ detection capabilities regardless of training modality.
With respect to hypothesis novelty and prior testing, H1 has been tested in small academic settings [39] but never at enterprise scale; H2 and H3 build on established training effectiveness research [22, 7, 21] but test these findings in a large-scale operational context using standardized difficulty measures; and H4 is novel, addressing the research gap identified by Franz et al. [10] regarding interactions between training modality and phishing complexity, which has received minimal empirical attention in prior work.
@@ 3.2 Study Design
To test the hypotheses, we employed a between-subjects experimental design [35]. Subjects were randomly assigned to one of three experimental conditions using our organization’s training management system. Each condition reflected the phishing training they received (Appendix 3.3). Subjects were recruited (Appendix 3.4), trained, and later received simulated phishing emails (Appendix 3.5). We compared subjects’ performance in handling these phishing emails between the treatments using several standard metrics (Appendix 3.6).
3.3 Training Modalities
@@
For the purposes of this research, the principal researcher collaborated with a cybersecurity training vendor who produces comprehensive security awareness training materials, phishing simulation exercises designed to augment those materials, and implementation assistance for organizational deployment. This partnership enabled access to professionally developed training content and established phishing templates while maintaining the experimental rigor necessary for our evaluation.
@@
The vendor offers three distinct training modalities, though we focused our evaluation on two of these approaches. The primary modality for cybersecurity awareness training consists of lecture-based videos that discuss the dangers of various social engineering attacks, including phishing recognition tactics, social engineering techniques, and organizational security protocols. Our vendor, along with many others in the industry, supplements this foundational training with interactive phishing exercises that present users with samples from a library of approximately 20 phishing templates. These exercises guide users through identifying suspicious elements and security deficiencies in emails that roughly correspond to the cues assessed by the NIST Phish Scale.
The vendor also provides a third modality referred to as “in-place training.” This approach occurs during phishing simulations when users click on malicious links, instead of being directed to a malicious website, they are immediately redirected to a training webpage that explains why the email they clicked should have been detected as fraudulent. However, decades of research into the effectiveness of embedded training environments [23] and recent comprehensive research by Ho et al. [4] convinced us that monitoring the results of in-place training was not necessary for our research objectives. Ho’s large-scale study across 19,500 healthcare employees found minimal effectiveness of embedded training approaches, providing thorough and convincing evidence that this modality offers limited practical value compared to the lecture-based and interactive approaches we chose to evaluate. We employed two distinct training approaches in collaboration with our cybersecurity training vendor, building on established taxonomies of phishing interventions [10]. Both modalities fall within Franz et al.’s “training” category [10], specifically comparing passive lecture-based delivery Figure 2against interactive approaches with embedded learning elements.
Treatment A: Lecture-Based Training: The lecture-style training consisted of fifteen brief instructional videos, each lasting 1-2 minutes, covering core cybersecurity awareness topics including phishing recognition tactics, social engineering techniques, and organizational security protocols [22]. The videos specifically addressed phishing identification strategies, emphasizing visual cues such as suspicious URLs, sender inconsistencies, and urgency tactics commonly employed by attackers. Following the video content, participants completed a knowledge assessment quiz reviewing key concepts from the training materials. This approach represents the traditional passive learning modality widely deployed in organizational cybersecurity training programs [23].
Figure 2: Traditional Training: Videos and Quiz
Treatment B: Interactive Training (Training + Exercise): The interactive group received the identical lecture-style training described above, plus an additional hands-on component consisting of simulated phishing exercises with immediate feedback. During these exercises, participants were presented with sample emails and asked to identify whether each message was legitimate or malicious. The system provided real-time feedback explaining the reasoning behind correct identification, highlighting specific phishing indicators, and reinforcing appropriate response behaviors such as using the “Report Phish” button in the email client [22]. This interactive component was designed to provide experiential learning opportunities that complement the passive video instruction, following research suggesting that active engagement enhances retention and behavioral change in cybersecurity training contexts [33].
Control Group: The control group received no cybersecurity training prior to the phishing simulation phase. This group was included to establish baseline susceptibility measurements. Control group participants received the same training materials after completion of the study to ensure organizational compliance with cybersecurity awareness requirements. Organizational constraints compelled us to use a relatively small control group, as ACME Corp. required that the majority of employees receive some form of cybersecurity training for compliance purposes.
3.4 Subject Recruitment
@@
Study Context: The subjects were employees at a U.S.-based international financial technology (“fintech”) firm. Participants represented a diverse cross-section of the organization, including employees from finance, technology, operations, customer service, and administrative functions. All participants were full-time employees with active email accounts who regularly use organizational communication systems and speak English.
@@
Subject Recruitment: This study leveraged organizational security training data collected as part of routine cybersecurity awareness training. The training was introduced through official organizational communications emphasizing its importance for maintaining security compliance and protecting organizational assets. This training was mandatory due to the organization’s need for regulatory compliance (cf. Appendix 2.1.2).
@@
Randomization: Subjects wre randomly assigned to each treatment. To mitigate bias by organizational role or function, we used stratified sampling [41] across departments to ensure representative distributions in each treatment group.
Completion Rates: Our data came from the 12,511 participants who completed the full experimental protocol. Table IV shows the completion rates by treatment type. The comparable completion rates across conditions (ranging from 87.3% to 89.1%) suggest that our random sample was preserved between recruitment and completion.
TABLE IV: Recruitment Pool and Completion Rates by Experimental Condition. Note the comparable drop-out rates by condition.
@@
Training Only (Treatment A) 6,758 6,023 (89.1%)
Training + Exercise (Treatment B) 6,764 6,026 (89.1%)
@@
Ethical Considerations: We acknowledge that employee consent in workplace security training cannot be considered fully voluntary due to inherent power dynamics and employment dependencies [5]. However, several ethical safeguards were incorporated: all employees received security training regardless of group assignment, individual performance data was anonymized and never used for personnel decisions, and phishing simulation templates excluded content designed to cause psychological distress (Appendix 3.5).
The research protocol was reviewed by the academic institution’s IRB, which determined that analysis of operational security training data does not constitute human subjects research under federal guidelines, as it involves analysis of data collected for legitimate business purposes rather than research-specific data collection [5].
See Appendix 3 for an extended discussion of ethics.
3.5 Phishing Simulations
After training, all subjects received one simulated phishing email, of varying complexity. We first describe how we defined complexity using NIST Phish Scale, and then how we deployed the phishing simulation.
3.5.1 Template Selection using NIST Phish Scale
To establish controlled and consistent evaluation of participant responses, we systematically assessed and categorized the difficulty of each simulated phishing email using a structured expert review process informed by the NIST Phish Scale framework [33]. Phishing templates were initially selected from our cybersecurity training vendor’s library, which provided preliminary difficulty assessments (easy, medium, hard). We then independently evaluated these templates using the official NIST Phish Scale User Guide methodology to ensure accurate difficulty classification.
The assessment team comprised three members of ACME Corp.’s IT security staff, each with substantial experience in identifying real-world phishing attacks and analyzing user behavior in response to simulated campaigns. This multi-rater approach follows established practices for enhancing reliability and validity of stimulus assessment in cybersecurity research [19]. Two team members independently evaluated each phishing email template, rating the perceived human difficulty in correctly identifying the email as malicious using structured criteria derived from the NIST Phish Scale methodology.
Our implementation followed the official NIST Phish Scale User Guide [33], evaluating templates across two primary dimensions. For the Phishing Cues dimension, raters assessed observable indicators that might alert users to fraudulent content, including spelling errors, suspicious URLs, formatting inconsistencies, and sender authenticity markers [34]. Fewer or less obvious cues indicated higher difficulty ratings.
For the Premise Alignment dimension, raters evaluated how well each email’s subject matter, sender, and content aligned with typical organizational communications that ACME Corp. employees receive in their regular workflow. This assessment considered whether the email appeared to originate from familiar organizational systems (e.g., HR portal, IT helpdesk, financial systems), addressed topics relevant to employee responsibilities, and used language and formatting consistent with legitimate internal communications [6]. Higher premise alignment—emails that closely resembled routine organizational correspondence—contributed to increased difficulty ratings.
Following independent assessment, the two raters compared their evaluations for each email template. Any disagreements regarding difficulty classification were discussed to identify the specific cues or contextual elements that led to differing conclusions. A third senior IT security expert then reviewed each email and the rationales provided by the independent raters, serving as a final validator to reach consensus difficulty ratings for each template. This collaborative validation process ensured that our phishing simulation represented a controlled range of difficulty levels essential for evaluating training effectiveness across standardized complexity measures [39].
Difficulty Categorization and Distribution: Following the NIST framework, we created a 3x3 classification matrix combining cue visibility (few-low, some-medium, many-high) with premise alignment (weak-low, medium, strong-high). Templates were assigned composite difficulty scores by summing both dimensions, with scores ranging from 2 (easy: high cues, low alignment) to 6 (hard: low cues, high alignment).
Our final simulation campaign included nine distinct phishing templates across three difficulty categories:
  • Easy: 5,721 emails (45.7% of total simulations)
  • Medium: 2,279 emails (18.2% of total simulations)
  • Hard: 4,511 emails (36.1% of total simulations)
(a) “Easy”: low premise, high cues.
(b) “Hard”: high premise, low cues.
Figure 3: Example templates. The “Easy” template has low premise alignment and high cues; the “Hard” is the inverse.
3.5.2 Simulation Deployment
Phishing simulations were distributed among participants within a three-month window after training completion. Each participant received one phishing email. Templates were randomly assigned to ensure balanced distribution across experimental conditions.
The deployed templates were constrained by organizational considerations. ACME Corp.’s management requested the exclusion of highly distressing content (e.g., emergency family situations, personnel security scenarios) to minimize employee psychological impact. This constraint reflects the practical limitations of conducting research within operational environments, where employee welfare considerations influence experimental designs [9].
Technical delivery failures, including bounced emails and system errors, resulted in participant exclusion from analysis to maintain data integrity.
@@ 3.6 Metrics
Our outcome measures followed established practices in phishing research [22, 23].
  • Open Rate captured the percentage of recipients who opened the simulated phishing email, measured via embedded tracking pixels, though we acknowledge this metric can be influenced by email client settings and security configurations [36].
  • Click-Through Rate (CTR) measured the percentage of recipients who clicked on malicious links within the phishing emails, serving as our primary indicator of phishing susceptibility [7].
  • Reporting Rate tracked the percentage of recipients who reported the suspicious email to the organization’s security team using the integrated “Report Phish” button, representing proactive security behavior and successful threat recognition [9].
The simulation platform tracked these metrics as follows: email open rates were tracked via embedded tracking pixels, click-through rates were tracked via unique URLs, and reporting rates were tracked via integration with the organization’s security incident reporting system.
@@ 3.7 Analysis
Data Validation and Quality Assurance: Prior to analysis, we conducted extensive data validation to ensure analytical integrity. This included verification of random assignment effectiveness across groups, assessment of technical delivery success rates, and examination of potential outliers or data anomalies. Participants with incomplete data due to technical failures, extended leave, or incomplete training were excluded from analysis to maintain dataset quality [39].
Statistical Approach: We employed two complementary analytical approaches to examine training effectiveness. Our primary analysis used two-way analysis of variance (ANOVA) to test main effects and interactions between training modality (Control, Training Only, Training + Exercise) and phishing difficulty level (Easy, Medium, Hard) on our outcome measures. Note that although ANOVA with binary outcomes violates standard normality assumptions, this approach aligns with established practices in phishing research [22, 23] and enables direct comparison with prior studies.
To address the limitations of ANOVA with binary data, we conducted confirmatory analyses using logistic regression models appropriate for binary outcomes (clicked/didn’t click, reported/didn’t report). The logistic regression models employed maximum likelihood estimation and provided odds ratios for effect size interpretation. Both analytical approaches yielded consistent conclusions regarding training effectiveness, with ANOVA providing F-statistics for comparison with prior literature and logistic regression offering more statistically appropriate modeling of our binary outcomes. We report both approaches to provide comprehensive statistical evidence while acknowledging the methodological trade-offs inherent in each approach. All analyses were conducted using Python 3.9 with statsmodels and scikit-learn libraries for statistical modeling and cross-validation.

### v2

@@ 3 Methodology
To summarize the preceding analysis: phishing training is widely practiced, though recent evidence sheds doubt on its utility. Few studies have rigorously assessed outcomes against standardized difficulty measures at sufficient scale for robust inference. We provide a reproduction of this observation, tested through hypotheses derived from prior works.
In this section, we describe our methodology: study design, subjects, hypotheses, metrics, and analysis methods. For ethics considerations, see the associated appendix.
@@ 3.1 Study Design
Figure 2: Illustration of our methodology. Our two-factor design randomly assigns different trainings to different subjects, and then randomly sends phishing lures of varying complexity to the subjects. We compared between-subject performance using a mix of standard and novel metrics to test our five hypotheses.
@@
Our study design is illustrated in Figure˜2. We employed a two-factor between-subjects experimental design [47]. Subjects were randomly assigned to one of three experimental conditions using our organization’s training management system. Each condition reflected the phishing training they received. Subjects were randomly selected, trained, and later received simulated phishing emails of varying complexity. We quantitatively compared subjects’ performance in handling these phishing emails between the treatments.
@@
Factor 1—Training: The research team collaborated with a cybersecurity training vendor who produces comprehensive security awareness training materials, phishing simulation exercises designed to augment those materials, and implementation assistance for organizational deployment. This partnership enabled access to professional training content and established phishing templates.
@@
The vendor offers two distinct training approaches for phishing. The first approach, Treatment A, consists of lecture-based videos on the dangers of various social engineering attacks, including phishing recognition tactics, social engineering techniques, and organizational security protocols. Figure˜3 gives a screenshot. Comprehension is assessed through a quiz. The second approach, Treatment B, consists of those videos followed by a series of interactive phishing exercises that present users with samples from a library of approximately 20 phishing templates. These exercises guide users through identifying suspicious elements and security deficiencies in emails that roughly correspond to the cues assessed by the NIST Phish Scale, and reinforces appropriate response behaviors such as the “Report Phish” button in the email client [19]. Finally, we included a Control Group whose members received no cybersecurity training prior to the phishing simulation phase. This group was included to establish baseline susceptibility measurements. [^3] Control group participants received the same training materials after completion of the study to ensure organizational compliance with all cybersecurity awareness requirements. The study timing dictated a small control group, as ACME Corp. required that the majority of employees receive annual cybersecurity training for compliance purposes.
Figure 3: Training for Treatment A: Videos and Quiz
Factor 2—Phishing Lure Complexity: After training, all subjects received one simulated phishing email, of varying complexity. To establish controlled and consistent evaluation of participant responses, we systematically assessed and categorized the difficulty of each simulated phishing email using a structured expert review process informed by the NIST Phish Scale framework [27].
Phishing templates were selected from the vendor’s library. We evaluated these templates using the official NIST Phish Scale User Guide methodology to ensure accurate difficulty classification. For the Phishing Cues dimension, raters assessed observable indicators that might alert users to fraudulent content, including spelling errors, suspicious URLs, formatting inconsistencies, and sender authenticity markers [16]. For the Premise Alignment dimension, raters evaluated how well each email’s subject matter, sender, and content aligned with typical organizational communications that ACME Corp. employees receive in their regular workflow [2].
The assessment team comprised three members of ACME Corp.’s IT security staff, each experienced in identifying real-world phishing attacks. Following Downs et al.’s advice, two team members independently evaluated each phishing email template [30]. Any disagreements regarding difficulty classification were discussed and resolved, incorporating a third rater as needed. This collaborative validation process ensured that our phishing simulation represented a controlled range of difficulty levels essential for evaluating training effectiveness across standardized complexity measures [8].
Our final simulation campaign used nineteen phishing templates that we divided into Easy (high cues/low alignment — 5721 emails), Medium (some cues, some alignment — 2279 emails) and Hard (few cues/high alignment — 4511 emails). Examples are given in Figure˜4. A full summary is given in Table˜6 in the Appendix.
(a) Easy: low premise, high cues.
(b) Hard: high premise, low cues.
Figure 4: Two Example Phishing Templates.
Simulation Deployment: Phishing simulations were distributed among participants within a three-month window after training completion. Each participant received one phishing email. Templates were randomly assigned to ensure balanced distribution across training conditions.
3.2 Subjects
@@
Study Context: The subjects were employees at a U.S.-based international financial technology (“fintech”) firm. Participants represented a diverse cross-section of the organization, including employees from finance, technology, operations, customer service, and administrative functions. All participants were full-time employees with active email accounts who regularly use organizational communication systems.
@@
Subject Recruitment: This study leveraged organizational security training data collected as part of routine cybersecurity awareness training. The training was introduced through official organizational communications emphasizing its importance for maintaining security compliance and protecting organizational assets. This training was mandatory due to the organization’s need for regulatory compliance (cf. Appendix˜2.1.3).
@@
Randomization: Subjects were randomly assigned to each treatment. To mitigate bias by organizational role or function, we used stratified sampling [91] across departments to ensure representative distributions in each treatment group.
Completion Rates: Our data came from the 12,511 participants who completed the full experimental protocol. Table˜3 shows the completion rates by treatment type. The comparable completion rates across conditions (ranging from 87.3% to 89.1%) suggest that our random sample was preserved between recruitment and completion.
Table 3: Recruitment pool and completion rates by condition. Note the comparable drop-out rates by condition.
@@
Treatment A (Lecture) 6,758 6,023 (89.1%)
Treatment B (Lecture + Exercises) 6,764 6,026 (89.1%)
@@
Table 4: Summary of study hypotheses, their theoretical grounding, and relation to prior work. The hypotheses address the effects of phishing email difficulty, training, and training modality on detection and reporting behaviors.
Hypothesis Theoretical and Empirical Basis Relation to Prior Work
H1 (Phish Scale Effect): Phishing emails with higher difficulty, as measured by the NIST Phish Scale, will have higher click-through rates than less difficult ones. Grounded in the foundational NIST Phish Scale framework [27] and validation studies [8]. Supported by phishing cue research [16] and systematic email analysis [99]. The NIST Phish Scale has been validated in lab settings [8]. We provide the first large-scale enterprise validation.
H2 (Training Impact–General): Training will increase the detection and reporting of phishing emails. Evidence of training efficacy, reduced click rates [19], improved detection accuracy [90], and short-term benefits [56]. Disputed [48]. Tests whether training effects from small-scale studies [19, 90] replicate at scale, addressing Franz et al.’s call [37].
H3 (Interactive Training Effect–Reporting): Interactive components yield higher reporting rates than lecture-only. From Franz et al.’s modality taxonomy [37], and evidence of the effectiveness of active and feedback-rich training [19, 56, 82]. Builds on academic findings of interactive training benefits [19, 56]. We provide the first test at organizational scale.
H4 (Interaction Effect): There will be an interaction between training modality and phishing difficulty, with interactive training more beneficial for easier lures. Emerges from cognitive load and decision strategy research [30], difficulty-aware detection patterns [99], and findings on human limitations [2]. Addresses the gap identified by Franz et al. [37] regarding interaction effects between training and phishing complexity, largely unexplored empirically.
H5 (Reporting Timeliness): Training will improve the timeliness of phishing reporting, with trained users reporting suspicious emails more quickly than untrained users. Grounded in studies showing that training enhances threat recognition speed [19]. Supported by evidence that security education improves decision-making efficiency [90]. While individual reporting speed has been measured [53], no studies have examined training effects on organizational-level reporting timeliness [88].
3.3 Hypotheses
Based on prior work on phishing training efficacy [19, 18, 44, 90, 37, 48] and documented differences in risk perception [2, 29], we formulated and tested five hypotheses. Table˜4 states each hypothesis, its basis, and its relation to prior work.
Our specific hypotheses are:
  1. 1.
    H1 (Phish Scale Effect): Phishing emails with higher difficulty, as measured by the NIST Phish Scale, will have higher click-through rates than less difficult ones.
  2. 2.
    H2 (Training Impact–General): Training will increase the detection and reporting of phishing emails.
  3. 3.
    H3 (Interactive Training Effect–Reporting): Interactive components yield higher reporting rates than lecture-only training.
  4. 4.
    H4 (Interaction Effect): There will be an interaction between training modality and phishing difficulty, with interactive training more beneficial for easier lures.
  5. 5.
    H5 (Reporting Timeliness): Training will improve the timeliness of phishing reporting, with trained users reporting suspicious emails more quickly than untrained users.
With respect to hypothesis novelty and prior testing: H1 has been tested in small academic settings [8] but never at enterprise scale. H2 and H3 build on established training effectiveness research [19, 90, 56] but applies these findings in a large-scale operational context using a standardized difficulty measure; H4 is novel, addressing the research gap identified by Franz et al. [37] regarding interactions between training modality and phishing complexity, which has received minimal empirical attention in prior work. H5 is also novel, directly addressing Steves et al.’s research question about organizational temporal patterns [88].
@@ 3.4 Metrics
Our outcome measures were a mix of standard metrics for comparison to prior results [19, 18], and novel metrics related to response timeliness per Steves et al. [88]. We performed measurements using the vendor’s simulation platform, which reported the standard metrics directly and timing information that we used to derive the novel metrics.
Standard Metrics: We employed three standard behavioral metrics to assess training effectiveness [19, 18]:
  • Open Rate is the percent of recipients who open the simulated phishing email, measured via embedded tracking pixels. [^4] This is the standard definition, though we acknowledge it can be influenced by email client settings and security configurations [6].
  • Click-Through Rate (CTR) is the percent of recipients who click on malicious links within the phishing emails. CTR is our primary indicator of phishing susceptibility [90].
  • Reporting Rate is the percent of recipients who report the suspicious email to the organization’s security team using the integrated “Report Phish” button, representing proactive security behavior and successful threat recognition [29].
Novel Temporal Metrics: Abroshan et al. suggested examining temporal patterns in user responses [2], and Steves et al. asked: “Is time to first report sooner than time to first click?” [88]. Rapid reporting of phish attempts enables the feedback loop built into modern email defenses (Figure˜1). We introduce the following metrics to see organizational-level temporal security dynamics:
  • Individual Reporting Timeliness measures how quickly individual users report suspicious emails:
    $\text{Timeliness}_{i}=t_{\text{report},i}-t_{\text{deployment}}$
    where $t_{\text{report},i}$ is the timestamp when user $i$ reported the email and $t_{\text{deployment}}$ is when the phishing campaign was launched.
  • Campaign Inoculation Status determines whether an organization can mitigate a phishing campaign before exploitation occurs:
    $\text{Inoculation}=\begin{cases}1&\text{if }t_{\text{first report}}<t_{\text{first click}}\\ 0&\text{otherwise}\end{cases}$
    where $t_{\text{first report}}$ and $t_{\text{first click}}$ are the timestamps of the first report and first click for a given campaign, respectively.
  • Organizational Inoculation Index quantifies the temporal advantage (or disadvantage) of reporting over clicking:
    $\text{OII}=t_{\text{first click}}-t_{\text{first report}}$
    Positive values indicate “inoculation” where reports preceded clicks, enabling organizational threat mitigation before exploitation. Negative values indicate that clicks occurred before reports, suggesting vulnerability windows.
This novel organizational-level metric directly addresses Steves et al.’s research question while providing actionable insights into collective security behaviors that complement individual-focused assessments. The Organizational Inoculation Index enables organizations to evaluate whether their workforce’s collective reporting behavior provides sufficient early warning to enable proactive threat response.
@@ 3.5 Analysis
Data Validation and Quality Assurance: Prior to analysis, we conducted data validation to ensure analytical integrity. This included verification of random assignment effectiveness across groups, assessment of technical delivery success rates, and examination of potential outliers or data anomalies. Participants with incomplete data due to technical failures, extended leave, or incomplete training were excluded from analysis to maintain dataset quality [8].
Statistical Approach: We employed two complementary analytical approaches to examine training effectiveness. Our primary analysis used two-way analysis of variance (ANOVA) to test main effects and interactions between training modality (Treatments A, B, and Control) and phishing difficulty level (Easy, Medium, Hard) on our outcome measures. Note that although ANOVA with binary outcomes violates standard normality assumptions, this approach aligns with established practices in phishing research [19, 18] and enables direct comparison with prior studies.
To address the limitations of ANOVA with binary data, we conducted confirmatory analyses using logistic regression models appropriate for binary outcomes (clicked/not, reported/not) [83]. The logistic regression models employed maximum likelihood estimation and provided odds ratios for effect size interpretation. Both analytical approaches yielded consistent conclusions regarding training effectiveness, with ANOVA providing F-statistics for comparison with prior literature and logistic regression offering more statistically appropriate modeling of our binary outcomes. We report both approaches to provide comprehensive statistical evidence while acknowledging the methodological trade-offs inherent in each approach. All analyses were conducted using Python [77] with pandas [57] for data manipulation, NumPy [45] for numerical computations, SciPy [95] for statistical tests, and statsmodels [83] for regression modeling and maximum likelihood estimation. Visualizations were created using matplotlib [50] and seaborn [97].